One pipe fills a pool in 6 hours. Second pipe can fill the same pool in 4 hours. How long would it take to fill the pool if both pipes were working.
Answer by stanbon(75887)

One pipe fills a pool in 6 hours. Second pipe can fill the same pool in 4 hours. How long would it take to fill the pool if both pipes were working.
------
1st pipe rate = 1/6 job/hr
2nd pipe rate = 1/4 job/hr
Together rate = 1/x job/hr
--------------------------------
Equation:
rate + rate = together rate
1/6 + 1/4 = 1/x
2x + 3x = 12
5x = 12
x = 12/5 hrs = 2.4 hrs (time to fill the pool if both pipes are working)
=======================
Cheers,
Stan H.